# Towards Scalable Automated Alignment of LLMs: A Survey

### v1

@@ 4.2.1 Response-Guided Distillation
In response-guided distillation, the target model emulates the teacher model by directly learning the responses to different instructions through instruction tuning. This approach has inspired numerous studies that aim to distill various capabilities from the teacher model to the target model. These capabilities include not only general instruction-following skills but also domain-specific abilities such as mathematics, coding, and agent-related tasks
@@
After constructing instruction data, corresponding responses can easily be developed from the teacher model. Training with these instruction-response pairs emulates the teacher’s capability of following instructions. For instance, LLaMA-GPT4 (Peng et al., 2023) utilizes GPT-4 to generate responses to instructions derived from Alpaca (Taori et al., 2023). Ubani et al. (2023) discuss the zero-shot prompting of ChatGPT for data augmentation in low-resource scenarios, synthesizing instructional data by designing high-quality and diverse prompts. More related works can be found in the last subsection. In addition to single-round data, some studies focus on collecting multi-turn trajectories from teacher models. Baize (Xu et al., 2023b) and Ultrachat (Ding et al., 2023) employ a comprehensive framework, using two ChatGPT APIs to play the roles of user and assistant to generate multi-round conversations. Parrot (Sun et al., 2023b) trains models to simulate humans in generating instructions and uses these trained models to engage in multi-turn conversations with ChatGPT on various topics, resulting in a collection of 40,000 multi-turn dialogues.
@@
Wizardmath (Luo et al., 2023a) employs the Evol-Instruct method to construct a comprehensive dataset specifically for mathematical reasoning tasks. MetaMath (Yu et al., 2024b) utilizes ChatGPT to bootstrap mathematical questions by rephrasing them from multiple perspectives without introducing additional knowledge. MAmmoTH (Yue et al., 2024a) produces a dataset comprising math problems and model-generated solutions, distinguished by a unique combination of chain-of-thought (CoT) and program-of-thought (PoT) rationales. MathCoder (Wang et al., 2024d) generates innovative and high-quality math problems along with their code-based solutions using the GPT-4 Code Interpreter. MathGenie (Lu et al., 2024) generates diverse and reliable math problems through a process of question back-translation. MARIO (Liao et al., 2024) leverages GSM8K and MATH as seed data, resulting in a collection of 26.9K solutions annotated by both GPT and human experts. Beyond purely mathematical data, several other studies propose transferring essential reasoning abilities from commercial LLMs to small models by generating detailed COT responses (Shridhar et al., 2023; Fu et al., 2023b; Hsieh et al., 2023; Magister et al., 2023; Ho et al., 2023; Li et al., 2022, 2023a).
@@
State-of-the-art LLMs, such as GPT-4, demonstrate exceptional performance in coding tasks. Apart from pre-training on raw code data, some approaches aim to transfer coding capabilities from teacher models through instruction tuning. Code Alpaca (Chaudhary, 2023) and WizardCoder (Luo et al., 2024) adhere to general automatic instruction-building paradigms. Code Alpaca employs Self-Instruct on 20K instruction-following data, thereby extending Alpaca’s capabilities to the coding domain. WizardCoder adapts the Evol-Instruct method for the coding domain, generating complex code and program instructions from simple coding and programming directives. WaveCoder (Yu et al., 2024c) and Magicoder (Wei et al., 2023) create high-quality instruction data utilizing open-source code datasets. WaveCoder enhances LLMs with open-source code snippets to produce superior instruction data for coding tasks. Magicoder creates multi-task data generated according to the techniques of the Self-Instruct. OpenCodeInterpreter (Zheng et al., 2024c) utilizes GPT-3.5 and GPT-4 to improve solutions with integrated text explanations and code snippets, incorporating execution and feedback for dynamic code refinement.
@@
Although open-source LLMs have achieved comparable performance to commercial models in many aspects, their capabilities in agent-related functions, such as tool usage and complex task planning, remain significantly limited. To address this issue, ToolLLM (Qin et al., 2023) has created an instruction-tuning dataset named ToolBench with ChatGPT, acquiring general tool usage capabilities in a zero-shot manner. Similar works include Graph-ToolFormer (Zhang, 2023), Gorilla (Patil et al., 2023), GPT4Tools (Yang et al., 2023b), ToolAlpaca (Tang et al., 2023), and others. Beyond tool usage, some studies focus on planning tasks. Examples include FIREACT (Chen et al., 2023a), AgentTuning (Zeng et al., 2023), ReAct Meets ActRe (Aksitov et al., 2023), ReST meets ReAct (Yang et al., 2024c), and ETO (Song et al., 2024b).
@@ 4.2.2 Preference-Guided Distillation
Although response-guided distillation can enhance the performance of student models (Wang et al., 2022), it does not effectively help the student model align with human preferences (Xu et al., 2024c). Therefore, some works concentrate on preference-guided distillation, which aligns the student model with the preferences reflected in the output from the teacher model. In this paradigm, the teacher model is guided to generate preference data in the form of partial order pairs, which are then used to align the student model via direct preference optimization algorithms such as DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), and PRO (Song et al., 2023). Based on the methodologies for constructing partial order signals, current works primarily encompass three paradigms: 1) Score-based:, which involves scoring and ranking responses; 2) Refine-based, which involves refining existing responses with AI feedback; and 3) Source-based, which focuses on learning the human preference of different data sources.
@@
Other studies improve initial responses using powerful models. Aligner (Ji et al., 2024) and MetaAligner (Yang et al., 2024a) utilize models such as GPT-4 to revise original responses and construct preference data. IterAlign (Chen et al., 2024e) automatically discovers new constitutions using an LLM and optimizes responses generated from a red team dataset to create preference data. Safer-Instruct (Shi et al., 2024) employs reversed instruction tuning, instruction induction, and expert model evaluation, using both raw text and GPT-4 generated responses to build high-quality preference data.
@@ 4.3 Weak-to-Strong Alignment
As we mentioned in Section 1, the challenges of scalable oversight become a significant barrier for the continuous development of AI systems. Specifically, the difficulty lies in effectively providing supervision as the capabilities of AI systems gradually surpass those of humans. Given the impracticality of strong-to-weak distillation approaches, weak-to-strong alignment has emerged as one of the most promising directions for achieving automated scalable oversight (Burns et al., 2023). Previous studies have mainly focused on weak-to-strong generalization between humans and AI, such as the Iterated Amplification method (Christiano et al., 2018), which supervises strong learners by iteratively amplifying weak experts. Recent research has begun to explore using weaker models to guide stronger models to achieve superalignment (Christiano et al., 2018; Burns et al., 2023; Liu and Alahi, 2024). Based on the source of the alignment signals, these works can be categorized into two types: 1) using smaller but aligned models to generate signals, and 2) using weaker models to guide stronger models in generating signals. Moreover, some studies investigate whether models can learn from behaviors in easy tasks to improve their performance in more challenging tasks, which, although not classic behavior imitation, is still noteworthy (Hase et al., 2024; Sun et al., 2024b). In the following subsections, we will introduce the representative studies in each category respectively.
Burns et al. (2023) employ a weaker LLM as the teacher to train a stronger LLM using a weak-to-strong approach. They fine-tune a larger pre-trained model based on labels generated by the smaller, but aligned, model and observe that the larger target model consistently outperforms the smaller supervisory model. Instead of relying on a single teacher, Liu and Alahi (2024) aim to further enhance the alignment of strong models by co-supervising a powerful student with a diverse group of professional teachers. Somerstep et al. (2024) examine weak-to-strong generalization as a transfer learning problem, achieving this through a label refinement procedure. Additionally, Aligner (Ji et al., 2024) and MetaAligner (Yang et al., 2024a) create partial order data by using a significantly smaller but aligned model to optimize the responses from stronger models.
@@ 4.4 Discussion
Achieving superalignment remains a significant challenge. We still lack a strong scientific understanding of superalignment (Burns et al., 2023), hindering further exploration of weak-to-strong alignment. Additionally, most current approaches still require a sufficiently aligned “weak” model, and how to utilize a truly weak model for superalignment remains an issue. Some works present theoretical frameworks for understanding weak-to-strong generalization (Charikar et al., 2024; Lang et al., 2024; Somerstep et al., 2024), but still have limited application scope. An interesting road is like ExPO (Zheng et al., 2024a), it extrapolates directly from an SFT model and an aligned model’s weights, obtaining a better-aligned model without additional training, demonstrating a promising approach from weak to strong.
@@ 5 Aligning Through Model Feedback
Human feedback reflects human values and can be used to align the LLMs, enabling LLMs to produce helpful and safe responses while correcting errors and toxic outputs. Unfortunately, accessing human feedback during training is challenging due to inefficiency and high costs. To address this issue, model feedback is introduced as a way to estimate human feedback. This approach is commonly utilized in reinforcement learning, where a reward model generates feedback. Compared to relying on the limited feedback data generated by humans, the reward model can perform feedback prediction across a broader distribution, thus achieving more efficient alignment. Aligning through automated generated model feedback offers an effective method for aligning LLM with human values, presenting a promising path towards achieving automated alignment. In this section, we explain how to leverage model-provided feedback to align it with human values. As shown in Figure 7, the methods can be divided into three types based on the form of feedback signals: scalar (§ 5.1), binary (§ 5.2), and text signals (§ 5.3).
Figure 7: The illustration of aligning through model feedback generated by reward model. The reward model will automatically generate the feedback of LLMs’ responses in the scalar, binary or text format.
@@ 5.1 Scalar Reward
Scalar signals are commonly generated by a reward model that takes the response of LLMs as input to generate scalar signals for estimating human preferences. Reward model is frequently employed in reinforcement learning to align LLM with human values. In this way, LLMs can automatically align with human values by utilizing the large amount and diverse feedback provided by the reward model. To achieve more effective automated alignment, recent studies focus on how to train a higher-quality reward model, and reducing the reliance on human annotation during the training of the reward model through model generation or pre-training. Besides, the scalar signals generated by reward model can also be used to optimize the generation of LLM during decoding and filter training data for instruct-tuning.
@@ 5.1.2 Improvement for Reward Modeling
To achieve more effective automated alignment, it is crucial to improve the quality of model feedback. Therefore, recent studies focus on learning high-quality reward models. The primary challenges in training reward model involve data collection and model optimization. Preference data collected is typically sparse and deficient in consistency and detail, and model optimization can be hindered by issues such as over-fitting.
@@
Because human annotators have different evaluation principles and subjective perspectives, the feedback is diverse and includes multiple viewpoints. Previous studies have used strategies like multiply models ensemble (Rame et al., 2023; Touvron et al., 2023), multi-objective learning (Zeng et al., 2024) to mitigate the negativity from the diversity data. In contrast to reward models that produce a single score,  Li et al. (2024a) introduce the Distributional Preference Reward Model (DPRM) for predicting preference distributions.
@@
The learning process of the reward model usually faces the problem of over-optimization. That is, through learning, the reward model performs poorly instead.  Gao et al. (2023) analyze this phenomenon through experiments and discovered the scaling law of the reward model to guide learning. Zhu et al. (2023a) provide a theoretical analysis of the reward model training in RLHF and show the importance of introducing pessimism when training. Moreover, some other works have adopted various ways to improve the performance of the reward model, including normalization (Zheng et al., 2023) and iterative learning (Touvron et al., 2023).
Although the purpose of the reward model is to predict human feedback, the reward modeling is challenging. Therefore, how to build a more comprehensive reward model to achieve automated alignment is an important research question.
@@ 5.1.3 Reinforcement Learning from AI Feedback
Reward model is commonly trained using human feedback which is hard and expensive to annotate. For the purpose of reducing human effort and improving the automation in alignment, there are some works that use the existing large language models to generate the preference data. Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023) trains the reward model with preference data of LLM and it can achieve comparable or superior performance to RLHF. The methods are mainly divided into two types, including ranking the responses of multiple models and directly generating positive and negative responses. In this way, the automated alignment can be achieved in the entire procedure of reinforcement learning without human effort.
@@
With the improved capabilities of LLMs, directly using them to rank multiple responses can provide preference data (Tunstall et al., 2023; Hong et al., 2023; Guo et al., 2024b; Pace et al., 2024; Yuan et al., 2024). This ranked preference data can also be produced with minimal human supervision, such as through human-defined principles (Bai et al., 2022b; Sun et al., 2023c) or rules (Kim et al., 2023). To improve the quality of generated preference data, Shi et al. (2024) propose a carefully designed pipeline including reversed instruction tuning, instruction induction, and expert model evaluation.  Liu et al. (2024a) propose to score the response using contrastive prompt pairs as input, which can achieve better performance compared to directly generated feedback with single prompt.
@@
Some of works use LLM to generate preference data directly by prompting it to output positive response and negative response (Chen et al., 2024c).  Yang et al. (2024b) use an indirect approach by using different prompts to generate positive and negative responses respectively.

### v2

@@ 4.2.1 Response-Guided Distillation
In response-guided distillation, the target model emulates the teacher model by directly learning the responses to different instructions through instruction tuning. This approach has inspired numerous studies that aim to distill various capabilities from the teacher model to the target model. These capabilities include not only general instruction-following skills but also domain-specific abilities such as mathematics, coding, and agent-related tasks.
@@
After constructing instruction data, corresponding responses can easily be developed from the teacher model. Training with these instruction-response pairs emulates the teacher’s capability of following instructions. For instance, LLaMA-GPT4 (Peng et al., 2023) utilizes GPT-4 to generate responses to instructions derived from Alpaca (Taori et al., 2023). In addition to single-round data, some studies focus on collecting multi-turn trajectories from teacher models. Baize (Xu et al., 2023b) and Ultrachat (Ding et al., 2023) using two ChatGPT APIs to play the roles of user and assistant to generate multi-round conversations. Parrot (Sun et al., 2023b) trains models to simulate humans in generating instructions and uses these trained models to engage in multi-turn conversations with ChatGPT on various topics.
@@
Wizardmath (Luo et al., 2023a) employs the Evol-Instruct method to construct a comprehensive dataset specifically for mathematical reasoning tasks. MetaMath (Yu et al., 2024b) utilizes ChatGPT to bootstrap mathematical questions by rephrasing them from multiple perspectives without introducing additional knowledge. MAmmoTH (Yue et al., 2024a) produces a dataset comprising math problems and model-generated solutions distinguished by a unique combination of chain-of-thought (CoT) and program-of-thought (PoT) rationales. MathCoder (Wang et al., 2024d) generates innovative and high-quality math problems along with their code-based solutions using the GPT-4 Code Interpreter. MathGenie (Lu et al., 2024) generates diverse and reliable math problems through a process of question back-translation. MARIO (Liao et al., 2024) leverages GSM8K and MATH as seed data, resulting in 26.9K solutions annotated by GPT and human experts. Beyond purely mathematical data, several other studies propose transferring essential reasoning abilities from commercial LLMs to small models by generating detailed CoT responses (Shridhar et al., 2023; Fu et al., 2023b; Hsieh et al., 2023; Magister et al., 2023; Ho et al., 2023; Li et al., 2022, 2023a; Zhou et al., 2024; Hong et al., 2024).
@@
State-of-the-art LLMs, such as GPT-4, demonstrate exceptional performance in coding tasks. Apart from pre-training on raw code data, some approaches aim to transfer coding capabilities from teacher models through instruction tuning. Code Alpaca (Chaudhary, 2023) and WizardCoder (Luo et al., 2024) adhere to general automatic instruction-building paradigms. Code Alpaca employs Self-Instruct on 20K instruction-following data, thereby extending Alpaca’s capabilities to the coding domain. WizardCoder adapts the Evol-Instruct method for the coding domain, generating complex code and program instructions from simple coding and programming directives. WaveCoder (Yu et al., 2024c) and Magicoder (Wei et al., 2023) create high-quality instruction data utilizing open-source code datasets. WaveCoder enhances LLMs with open-source code snippets to produce superior instruction data for coding tasks. Magicoder creates multi-task data generated according to the techniques of the Self-Instruct. OpenCodeInterpreter (Zheng et al., 2024d) utilizes GPT-3.5 and GPT-4 to improve solutions with integrated text explanations and code snippets, incorporating execution and feedback for dynamic code refinement.
@@
Although open-source LLMs have achieved comparable performance to commercial models in many aspects, their capabilities in agent-related functions, such as tool usage and complex task planning, remain significantly limited. To address this issue, ToolLLM (Qin et al., 2023) has created an instruction-tuning dataset named ToolBench with ChatGPT, acquiring general tool usage capabilities in a zero-shot manner. Similar works include Graph-ToolFormer (Zhang, 2023), Gorilla (Patil et al., 2023), GPT4Tools (Yang et al., 2023b), ToolAlpaca (Tang et al., 2023), and others. Beyond tool usage, some studies focus on planning tasks. Examples include FIREACT (Chen et al., 2023a), AgentTuning (Zeng et al., 2023), ReAct Meets ActRe (Aksitov et al., 2023), ReST meets ReAct (Yang et al., 2024e), and ETO (Song et al., 2024b).
@@ 4.2.2 Preference-Guided Distillation
Although response-guided distillation can enhance the performance of student models (Wang et al., 2022), it does not effectively help the student model align with human preferences (Xu et al., 2024c). Therefore, some works concentrate on preference-guided distillation, which aligns the student model with the preferences reflected in the output from the teacher model. In this paradigm, the teacher model is guided to generate preference data in the form of partial order pairs, which are then used to align the student model via direct preference optimization algorithms such as DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), and PRO (Song et al., 2023). Based on the methodologies for constructing partial order signals, current works primarily encompass three paradigms: 1) Score-based, which involves scoring and ranking responses; 2) Refine-based, which involves refining existing responses with AI feedback; and 3) Source-based, which focuses on learning the human preference of different data sources.
@@
Other studies improve initial responses using powerful models. Aligner (Ji et al., 2024) and MetaAligner (Yang et al., 2024a) utilize models such as GPT-4 to revise original responses and construct preference data. IterAlign (Chen et al., 2024e) automatically discovers new constitutions using an LLM and optimizes responses generated from a red team dataset to create preference data. Safer-Instruct (Shi et al., 2024) employs reversed instruction tuning, instruction induction, and expert model evaluation, using both raw text and GPT-4 generated responses to build high-quality preference data. UltraInteract (Yuan et al., 2024a) build a preference tree for each instruction where trajectories are root-to-leaf paths and paired correct and incorrect nodes or trajectories can be used for alignment.
@@ 4.3 Weak-to-Strong Alignment
As we mentioned in Section 1, the challenges of scalable oversight become a significant barrier for the continuous development of AI systems. Specifically, the difficulty lies in effectively providing supervision as the capabilities of AI systems gradually surpass those of humans. Given the impracticality of strong-to-weak distillation approaches, weak-to-strong alignment has emerged as one of the most promising directions for achieving automated scalable oversight (Burns et al., 2023). Previous studies have mainly focused on weak-to-strong generalization between humans and AI, such as the Iterated Amplification method (Christiano et al., 2018), which supervises strong learners by iteratively amplifying weak experts. Recent research has begun to explore using weaker models to guide stronger models to achieve superalignment (Burns et al., 2023; Liu and Alahi, 2024). Based on the source of the alignment signals, these works can be categorized into two types: 1) using smaller but aligned models to generate signals, and 2) using weaker models to guide stronger models in generating signals. Moreover, some studies investigate whether models can learn from behaviors in easy tasks to improve their performance in more challenging tasks, which, although not classic behavior imitation, is still noteworthy (Hase et al., 2024; Sun et al., 2024b). In the following subsections, we will introduce the representative studies in each category respectively.
Burns et al. (2023) employ a weaker LLM as the teacher to train a stronger LLM using a weak-to-strong approach. They fine-tune a larger pre-trained model based on labels generated by the smaller but aligned model and observe that the larger target model consistently outperforms the smaller supervisory model. Instead of relying on a single teacher, Liu and Alahi (2024) aim to further enhance the alignment of strong models by co-supervising a powerful student with a diverse group of professional teachers. Somerstep et al. (2024) examine weak-to-strong generalization as a transfer learning problem, achieving this through a label refinement procedure. Yang et al. (2024d) study the multi-objective alignment in weak-to-strong generalization and discover that strong students may deceive weak teachers to gain high rewards in other dimensions, which can be mitigated by using an intermediate model. Additionally, Aligner (Ji et al., 2024) and MetaAligner (Yang et al., 2024a) create partial order data by using a significantly smaller but aligned model to optimize the responses from stronger models.
@@ 4.4 Discussion
Achieving superalignment remains a significant challenge. We still lack a strong scientific understanding of superalignment (Burns et al., 2023), hindering further exploration of weak-to-strong alignment. Additionally, most current approaches still require a sufficiently aligned ”weak” model, and how to utilize a truly weak model for superalignment remains an issue. Some works present theoretical frameworks for understanding weak-to-strong generalization (Charikar et al., 2024; Lang et al., 2024; Somerstep et al., 2024), but still have limited application scope. An interesting road is like ExPO (Zheng et al., 2024a). ExPO extrapolates directly from an SFT model and an aligned model’s weights, obtaining a better-aligned model without additional training, demonstrating a promising approach from weak to strong.
@@ 5 Aligning Through Model Feedback
Human feedback reflects human values and can be used to align the LLMs, enabling LLMs to produce helpful and safe responses while correcting errors and toxic outputs. Unfortunately, accessing human feedback during training is challenging due to inefficiency and high costs. To address this issue, model feedback is introduced as a way to estimate human feedback. This approach is commonly utilized in reinforcement learning, where a reward model generates feedback. Compared to relying on the limited feedback data generated by humans, the reward model can perform feedback prediction across a broader distribution, thus achieving more efficient alignment. Aligning through automated generated model feedback offers an effective method for aligning LLMs with human values, presenting a promising path towards achieving automated alignment. In this section, we explain how to leverage model-provided feedback to align it with human values. As shown in Figure 7, related methods can be divided into three types based on the form of feedback signals: scalar (§ 5.1), binary (§ 5.2), and text signals (§ 5.3).
Figure 7: The illustration of aligning through model feedback generated by reward models. The reward model will automatically generate the feedback of LLMs’ responses in the scalar, binary or text format.
@@ 5.1 Scalar Reward
Scalar signals are commonly generated by a reward model that takes the response of LLMs as input to generate scalar signals for estimating human preferences. Reward model is frequently employed in reinforcement learning to align LLM with human values. In this way, LLMs can automatically align with human values by utilizing the large amount and diverse feedback provided by the reward model. To achieve more effective automated alignment, recent studies focus on how to train a higher-quality reward model and reduce the reliance on human annotation during the training of the reward model through model generation or pre-training. Besides, the scalar signals generated by reward model can also be used to optimize the generation of LLM during decoding and filter training data for instruct-tuning.
@@ 5.1.2 Improvement for Reward Modeling
To achieve more effective automated alignment, it is crucial to improve the quality of model feedback. Therefore, recent studies focus on learning high-quality reward models. The primary challenges in training reward models involve data collection and model optimization. Collected Preference data is typically sparse and deficient in consistency and detail, and model optimization can be hindered by issues such as over-fitting.
@@
Because human annotators have different evaluation principles and subjective perspectives, the feedback is diverse and includes multiple viewpoints. Previous studies have used strategies like multiply models ensemble (Rame et al., 2023; Touvron et al., 2023), multi-objective learning (Zeng et al., 2024a; Zhong et al., 2024; Guo et al., 2024b; Yang et al., 2024c) to mitigate the negativity from the diversity data. In contrast to reward models that produce a single score,  Li et al. (2024a) introduce the Distributional Preference Reward Model (DPRM) for predicting preference distributions.
@@
The learning process of the reward model usually faces the problem of over-optimization. That is, through learning, the reward model performs poorly instead.  Gao et al. (2023) analyze this phenomenon through experiments and discover the scaling law of the reward model to guide learning. Zhu et al. (2023a) provide a theoretical analysis of the reward model training in RLHF and show the importance of introducing pessimism when training. Moreover, some other works have adopted various ways to improve the performance of the reward model, including normalization (Zheng et al., 2023) and iterative learning (Touvron et al., 2023).
Although the purpose of the reward model is to predict human feedback, modeling the reward is challenging. Therefore, how to build a more comprehensive reward model to achieve automated alignment is an important research question.
@@ 5.1.3 Reinforcement Learning from AI Feedback
Reward models are commonly trained using human feedback which is hard and expensive to annotate. For the purpose of reducing human effort and improving the automation in alignment, some works use the existing large language models to generate the preference data. Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023) trains the reward model with preference data of LLM and it can achieve comparable or superior performance to RLHF. The methods are mainly divided into two types, including ranking the responses of multiple models and directly generating positive and negative responses. In this way, the automated alignment can be achieved in the entire procedure of reinforcement learning without human effort.
@@
With the improved capabilities of LLMs, directly using them to rank multiple responses can provide preference data (Tunstall et al., 2023; Hong et al., 2023; Guo et al., 2024a; Pace et al., 2024; Yuan et al., 2024b). This ranked preference data can also be produced with minimal human supervision, such as through human-defined principles (Bai et al., 2022b; Sun et al., 2023c) or rules (Kim et al., 2023). To improve the quality of generated preference data, Shi et al. (2024) propose a carefully designed pipeline including reversed instruction tuning, instruction induction, and expert model evaluation.  Liu et al. (2024a) propose to score the response using contrastive prompt pairs as input, which can achieve better performance compared to directly generated feedback with a single prompt.
@@
Some works use LLM to generate preference data directly by prompting it to output positive response and negative response (Chen et al., 2024c).  Yang et al. (2024b) use an indirect approach by using different prompts to generate positive and negative responses, respectively.